Grade 11Mechanics

The length of the second hand of a clock is 1cm,what is the change in velocity of it`s tip in 15 seconds? Please answer as soon as possible and explain it. I have already seen the solution but I`m not able to understand why the negative sign is there with j cap. Please explain.

Adrita Das
8 Years agoGrade 11

2 Answers

Arun
8 Years ago

Since v(velocity) = r(radius)×w(angular veocity)

r= 1cm, w= (2pi/60) radian/sec, so v= (pi/30)cm/sec.

Now the sec hand has rotated by 90* and since the magnitude of the initial and final velocity of tip of sec hand is same. The change in magnitude can be calculated as V= sqrt(v sq +v sq)

So chnge in magnitude = sqrt2×(pi/30)

Khimraj
8 Years ago
magnitude of second hand is always constant only direction changes
After 15 sec change in direction will be 90 degrees
and constant velocity is given by
v = w.r = (2\pi/60)(0.01) m/s
SO change in velocity will be (0.01\pi/30)\sqrt{2} m/s
Hope it clears.
If you like answer then please approve it.